# Artificial Precision Timing Array: bridging the decihertz gravitational-wave sensitivity gap with clock satellites

### v1

@@ III Pulsar Timing Array Sensitivity
When collecting data, GW detectors are subject to several sources of noise that complicate the identification of signals from astrophysical sources. The data stream received by the detector $\mathbb{s}(t)$ is the sum of a noise term $\mathbb{n}(t)$ and a (possible) signal term $\mathbb{h}(t)$:
$\mathbb{s}(t)=\mathbb{n}(t)+\mathbb{h}(t).$ (4)
The significance of the signal is determined by the signal-to-noise ratio (SNR). Taking a frequentist approach to GW signal processing, one defines a statistic $\mathfrak{s}$ by cross-correlating the timing residuals with a symmetric filter matrix $\mathbb{K}(t)$:
$\mathfrak{s}=\int\mathrm{d}t\int\mathrm{d}t^{\prime}\mathbb{s}^{T}(t)\mathbb{K}^{\dagger}(t-t^{\prime})\mathbb{s}(t^{\prime})=\int\mathrm{d}f\;\mathbb{\tilde{s}}(f)^{T}\mathbb{\tilde{K}}^{\dagger}(f)\mathbb{\tilde{s}}(f),$ (5)
where the tilde denotes a Fourier transform. The SNR $\rho$ is the ratio of the average of this statistic in the presence of a signal to its root mean square (rms) value with only noise:
$\rho^{2}=\frac{\langle\mathfrak{s}\rangle^{2}_{s=n+h}}{\langle\mathfrak{s}^{2}\rangle_{s=n}-\langle\mathfrak{s}\rangle_{s=n}^{2}}.$ (6)
In GW signal processing, it is common practice to look for the “matched” filter $\mathbb{\tilde{K}}(f)$ that maximizes $\rho$ in order to optimally extract the signal from the noise. $\mathbb{\tilde{K}}(f)$ and its corresponding $\rho$ may be evaluated by assuming Gaussian white noise and that the pulsars are far enough from each other to ensure the noise of individual pulsars is uncorrelated Moore et al. (2015a). In this case, for a monochromatic GW emitted by a slowly evolving compact object binary, the SNR is
$\rho^{2}=\frac{1}{2}N_{p}(N_{p}-1)\frac{8\chi^{4}h_{c}^{4}}{T}\int_{1/T}^{1/\delta t}\mathrm{d}f\;\frac{\sin^{4}{\left[\pi(f-f_{0})T\right]}}{\pi^{4}(f-f_{0})^{4}f^{4}S_{n}^{2}},$ (7)
where $N_{p}$ is the number of pulsars, $1/\delta t$ is the cadence of observations, $T$ is the total time of observation of the source, and $S_{n}=2\sigma^{2}\delta t$ is the noise power spectral density for white noise, where $\sigma$ is the rms error in the timing residuals. $\chi$ is a sky-averaged version of the redshift $z$ of pulsar signals. In general, $z$ is
@@
$z=\frac{1}{2}\frac{\hat{p}^{i}\hat{p}^{j}}{1+\hat{\Omega}\cdot\hat{p}}\left(h_{ij}(t-|\mathbf{p}|(1+\hat{\Omega}\cdot\hat{p}),\hat{\Omega})-h_{ij}(t,\hat{\Omega})\right),$ (8)
where $t$ is the detection time, $\mathbf{p}$ is the position of the pulsar in an Earth-centered coordinate system, $\hat{\Omega}$ is the direction of propagation of the GW, and $h_{ij}$ is the GW perturbation to the background metric Anholm et al. (2009). In PTAs one often assumes that the pulsar term, i.e. the first term in Eq. 8, is zero Moore et al. (2015a). We discuss the implications of this in Section IV.2. Besides the sky-averaging, we average $z$ over GW polarization and source inclination angle so as not to overestimate the sensitivity of APTA, resulting in $\chi\approx 0.317$.
The integral in Eq. 7 can be analytically approximated in the high and low-frequency regimes Moore et al. (2015a). Choosing a detection-threshold value $\rho_{th}$ and inverting Eq. 7, the dimensionless characteristic GW strain $h_{c}$ in the high and low-frequency regimes is
$h_{c}^{\mathrm{HIGH}}(f)\approx\left(\frac{16\rho_{th}^{2}}{3\chi^{4}N_{p}(N_{p}-1)}\right)^{1/4}\sigma f\sqrt{\frac{\delta t}{T}}$ (9a)
$h_{c}^{\mathrm{LOW}}(f)\approx\frac{3\sqrt{\rho_{th}}}{2^{7/4}\chi\pi^{3}}\left(\frac{13}{N_{p}(N_{p}-1)}\right)^{1/4}\frac{\sigma}{T^{3}f^{2}}\sqrt{\frac{\delta t}{T}}\sec{\phi}$ (9b)
where $\phi$ is chosen to make $h_{c}^{\mathrm{HIGH}}(2/T)=h_{c}^{\mathrm{LOW}}(2/T)$. The total sensitivity is hence Moore et al. (2015a)
$h_{c}(f)\approx h_{c}^{\mathrm{HIGH}}(f)+h_{c}^{\mathrm{LOW}}(f).$ (10)
As this model shows, sensitivity curves fundamentally depend on the timing precision, the threshold SNR, the number of pulsars, and the cadence and total duration of observation.
@@ IV Artificial Precision Timing Array
To construct a GW detector based on PTA principles, we envision distributing pulsing electromagnetic sources in the sky. These light sources need to have their periodic pulses timed to high precision so that GWs may be observed with them. This can be achieved with an array of atomic-clock-carrying satellites, which would play the same role as the astrophysical pulsars in regular PTAs. They could orbit around the Earth or the Sun. Alternatively, rather than Earth-based radio receivers, one could imagine a near-Earth receiving base to avoid atmospheric interference, as illustrated in Fig. 1. The details of this communication mechanism, including the electromagnetic pulse features and its use at detection, are beyond the scope of this paper, which focuses on clock requirements. However, this communication pathway must not degrade the required clock accuracy.
@@
Figure 1: Cartoon illustration of the APTA concept. The detector is composed of an array of satellites communicating with a near-Earth central receiving station. The satellites, equipped with atomic clocks to allow for precision timing, play the role of the pulsars in this pulsar-timing-like GW detector. Objects are not drawn to scale, and the satellite array’s orbit may encircle some solar system planets, contrary to what the figure suggests.
@@
The results of the model from Section III do not depend on the pulsar/satellite communication frequency because $f_{photon}\gg f_{GW}$ for reasonable signal frequency choices Isaacson (1968); Stappers et al. (2011). To observe 0.1100.1100.1–100.1 – 10 Hz GWs, radio waves would do, but visible light would give a pulse-to-GW frequency ratio comparable to that of regular PTAs. X-rays may also become a good choice given recent advances in the generation of high-frequency phase-stable pulses (Zhang et al., 2022, e.g.). We leave the optimal photon wavelength for APTA pulses as a subject for future investigation.
IV.1 Timing error and cadence
@@
PTAs observing for a total time $T$ with a cadence of $1/\delta t$ are generally sensitive to GW frequencies in the range $(1/T)\lesssim f_{GW}\lesssim(1/\delta t)$ Moore et al. (2015b). Thus accessing the 0.1100.1100.1–100.1 – 10 Hz band can be achieved straightforwardly simply by searching for GW transients and continuous signals that last at least $\sim 10$ s and taking data at least every 0.1 s.
@@
The lowest detectable characteristic strain depends linearly on the error in the PTA timing residuals $\sigma$ as indicated in Eqs. III. We estimate the attainable sensitivities for APTA given various timing errors consistent with present and future atomic clock precision. We defer the discussion of additional sources of error to Section V. The best atomic clocks at present can reach fractional frequency uncertainties $\xi=\Delta f_{clock}/f_{clock}\sim 10^{-18}$ in 1 s of averaging and $\sim 10^{-19}–10^{-20}$ in 1 h Bothwell et al. (2022); Oelker et al. (2019); Campbell et al. (2017). In the next few years, we may reach $\xi\sim 10^{-20}$ in 1 s averaging with current 3D optical lattice clock technology. In a decade, use of squeezing and increased atom number may allow $\xi\sim 10^{-21}$ Ludlow et al. (2015). Note that $\xi\propto t_{avg}^{-1/2}$, for $t_{avg}$ the time of averaging. Our averaging time is limited by cadence so $\xi$ over time $\delta t$ of averaging is
@@
where $\xi_{1\text{ s}}$ is the fractional uncertainty in 1 s of averaging. We assume APTA’s clocks reach a stable optimal fractional uncertainty $\xi_{\delta t}$ after averging and use this value to compute the timing error as $\sigma=\xi_{\delta t}\delta t$.
@@
APTA can be sensitive to much higher frequencies than regular PTAs because its precision timing can be carried out continuously, whereas PTAs are limited by radio telescope access. Continued observation can be achieved if the beacon satellites are locked to a central observing station. In this case, the cadence $1/\delta t$ is bounded by the frequency at which communication pulses can be stably emitted and received, thereby fundamentally limited by the atomic clocks’ frequency.
For communication purposes, an extremely efficient way to transfer the timing precision of optical atomic clocks to the radiofrequency domain is provided by optical frequency comb technology Udem et al. (2002); Diddams et al. (2020). It has the potential to generate precision APTA pulses at a carrier frequency of 10 MHz to 100 GHz. We also note that intense efforts are underway to miniaturize and ruggedize the technology of strontium optical lattice clocks Poli et al. (2014); Koller et al. (2017); Origlia et al. (2018); Takamoto et al. (2020). Progress has been fast, which makes the prospect of deploying precision optical lattice clocks within the next decade highly realistic.
In the simplified timing model described in Section III, lowering $\delta t$ always lowers the sensitivity. This is a feature of assuming white noise: when $\delta t\rightarrow 0$, the measurement becomes continuous, so noise does not remain at any frequency. The effects of more realistic noise will be considered in future work. The current state of the art for phase-stable X-ray pulses is $10^{-15}$ s Zhang et al. (2022) For this study, we use $\delta t=0.01$ s, an order of magnitude below the minimum requirement for accessing 10 Hz GWs. Therefore, to focus on clock uncertainty requirements, we utilize state-of-the-art low values of $\xi$ alongside conservative values of $\delta t$. The same sensitivities could be reached with higher $\xi$ by lowering $\delta t$. The determination of the optimal set of parameters for constructing APTA is left for subsequent studies.
IV.2 Distance to the satellite array
The resulting sensitivity curve in Eqs. III and 10 does not depend on the distance between the pulsars and the observer because the correlated timing residual at each pulsar is assumed to be zero across all pulsars. Recent studies have considered the sensitivity enhancement from modeling non-zero pulsar terms in the timing residual (see discussion in Ref. Zhu et al. (2016) and references therein), but we leave a more detailed treatment of the pulsar term for future work. This approximation assumes that pulsars are far enough from each other that their separation is much larger than the GW wavelength (for stochastic GW searches) and consequently the GW phase is not coherent across the beacons. While this may not be the case for APTA in practice, we use this model for the purposes of this proof of concept study.
In this initial study, we assume operators have perfect control of the satellites. Hence, increasing the number of satellites $N_{s}$ will always increase the detector’s sensitivity. For a realistic scenario, we may consider a handful of satellites, $N_{s}\approx 6$ in orbit at roughly 1-2 AU from the sun. Although the assumption of uncorrelated satellite timing residuals may start to break down at these distances, the detector can still work with non-zero correlations between the satellite timing residuals included in the real sensitivity model.
IV.3 APTA sensitivity curves

### v2

@@ III.1 The Physics of Pulsar-Timing-Based GW Detection
We are concerned with first-order perturbations and the integrand $h^{TT}_{ij}$ is already a first-order term. Therefore, within the integral, we can consider the unperturbed trajectory of the photon,
$t_{o}$ $=t_{e}+d_{s}$ (8a)
$\vec{x_{0}}(t)$ $=(t_{e}-t+d_{s})\mathbf{\hat{p}}$, (8b)
such that
[…]
The fundamental idea in pulsar timing is to detect gravitational waves through deviations in the time elapsed between observed pulses. Therefore, consider the immediate next pulse, emitted by the satellite at time $t_{e}+\tau$, where $\tau$ is the pulsing period of the light source. The second pulse will be observed at $t_{o}^{\prime}$. The same derivation that led to Eq. 9 gives, by replacing $t_{o}$ and $t_{e}$ with $t_{o}^{\prime}$ and $t_{e}+\tau$,
[…]
Changing the integration variable in Eq. 10 as $t^{\prime}\rightarrow t^{\prime}-\tau$ and subtracting Eq. 9 from it, it becomes clear that the effect of the GW is a pulsing period shift as perceived by the observer:
$t_{o}^{\prime}-t_{o}=\tau+\Delta\tau$ (11a)
[…]
For the usual PTA, the typical source period $\tau_{p}$ is of order 1 ms while the target GW frequency $\omega_{gw,p}$ is of order $0.1\,\mathrm{yr^{-1}}$, such that $\omega_{gw,p}\tau_{p}\ll 1$. The validity of this assumption for APTA is discussed in Section IV.1. In this regime, we may Taylor expand $h^{TT}_{ij}(t^{\prime}+\tau,\vec{x_{0}}(t^{\prime}))$ around $t^{\prime}$ in powers of $\tau$ and discard second-order and higher terms:
$h^{TT}_{ij}(t^{\prime}+\tau,\vec{x_{0}}(t^{\prime}))=h_{ij}^{TT}(t^{\prime},\vec{x_{0}}(t^{\prime}))+\frac{\partial}{\partial t^{\prime}}h_{ij}^{TT}(t^{\prime},\vec{x_{0}}(t^{\prime}))\tau$ (12)
so that Eq. 11 becomes
[…]
To integrate Eq. 13, we now specialize to a monochromatic GW—which, depending on the observational timescale, may be a good approximation for several astrophysical sources—propagating along direction $\mathbf{\hat{n}}$ in TT gauge:
[…]
where the amplitude tensor satisfies the transverse condition $n^{i}\mathcal{A}_{ij}(\mathbf{\hat{n}})=0$. With this, Eq. 13 becomes:
[…]
PTA signal processing can be done with either the so-called redshift, expressed in Eq. 15, or with its integral, the timing residual $R(t)$:
[…]
where we have now reinserted factors of $c$ and set $t_{o}=t^{\prime}$ and $t_{e}=t^{\prime}-d_{s}/c$, and the constant term from integration was dropped. Ultimately, we will be interested in the frequency integral of $R(t)$’s Fourier transform, and that constant term will contribute a delta function centered at $f=0$ to that integrand, which will be outside of the range of integration in Eq. 22.
III.2 Gravitational-Wave-Detection Statistics
@@
When collecting data, GW detectors are subject to several sources of noise that complicate the identification of signals from astrophysical sources. The data stream $\mathbb{d}(t)$ received by the detector is the sum of a noise term $\mathbb{n}(t)$ and a (possible) signal term $\mathbb{h}(t)$. Consider $N$ detectors which have the data streams consisting of signals $h_{i}$ and additive noise $n_{i}$, $i=1,2,..,N$
$d_{i}(t)=h_{i}(t)+n_{i}(t).$ (17)
To detect the presence of a signal, we construct a test statistic (TS) from the likelihood ratio:
${\rm TS}=\frac{P(d|\mathcal{H}_{a})}{P(d|\mathcal{H}_{0})},$ (18)
where $\mathcal{H}_{a}$ and $\mathcal{H}_{0}$ denote the hypotheses where there is an astrophysical signal and no signal respectively. Assuming the noise is Gaussian and we are searching for a specific signal $\mathcal{T}_{a}$, the optimal detection statistic reduces to the well-known matched filter signal-to-noise ratio
[…]
where $\vec{x}_{i}$ are the detector positions.
When the sought astrophysical signal is detected ($D_{i}(f)=\mathcal{T}_{i}(f)$), $\rho$ is distributed approximately normally with variance 1 and mean
[…]
III.3 APTA Sensitivity Curve
With the tools built in Sections III.1 and III.2, we may now construct APTA’s sensitivity curve for a monochromatic GW, like what a slowly-evolving compact object binary emits.
Consider a source observed for a total time $T$ with observation cadence $1/\delta t$. Then, with the timing residual in Eq. 16 as the time-domain template, whose Fourier transform will be $\mathcal{T}_{i}(f)$ for the different satellites, and with white noise, for which $S_{1,n}=2\sigma^{2}\delta t$, where $\sigma$ is the rms error in the timing residuals, Eq. 21 becomes:
[…]
Without loss of generality, we can orient our coordinates such that $\mathbf{\hat{n}}=\mathbf{\hat{z}}=(0,0,1)$ and $\mathbf{\hat{p}}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta).$ Then, $\mathcal{A}_{ij}=h_{0}(A^{+}H_{ij}^{+}+A^{\times}H_{ij}^{\times})$, where $h_{0}$ is an overall amplitude determined by the astrophysical source; $A^{+}=(1+\cos^{2}\iota)/2$ and $A^{\times}=\cos\iota$, with $\iota$ being the inclination angle; and $H_{ij}^{+}=\epsilon_{ij}^{+}\cos(2\psi)+\epsilon_{ij}^{\times}\sin(2\psi)$ and $H_{ij}^{\times}=-\epsilon_{ij}^{+}\sin(2\psi)+\epsilon_{ij}^{\times}\cos(2\psi)$, with $\psi$ being the polarization angle, $\epsilon_{ij}^{+}=\hat{x}_{i}\hat{x}_{j}-\hat{y}_{i}\hat{y}_{j}$, and $\epsilon_{ij}^{\times}=\hat{x}_{i}\hat{y}_{j}+\hat{y}_{i}\hat{x}_{j}$ Moore et al. (2015a). Hence, the integrand in Eq. 22 becomes
@@ III.3 APTA Sensitivity Curve
$|\mathcal{T}_{i}(f)|^{2}\bigg{\rvert}_{f>0}=\frac{\delta_{T}(f-f_{\rm gw})^{2}}{8\pi^{2}f_{\rm gw}^{2}}\left(\frac{\hat{p}^{i}\hat{p}^{j}\mathcal{A}_{ij}}{2(1+\mathbf{\hat{p}}\cdot\mathbf{\hat{n}})}\right)^{2}\\ \times\left\{1-\cos\left[4\pi f_{\rm gw}\frac{d_{s}}{c}\cos^{2}\left(\frac{\theta}{2}\right)\right]\right\}.$ (23)
where $\delta_{T}(f)=\sin(\pi fT)/(\pi f)$ is the finite-time delta function. Instead of summing over satellites with different configurations, we determine the average satellite configuration by averaging $|\mathcal{T}_{i}(f)|^{2}$ over $\theta$, $\phi$, $\iota$, and $\psi$ and replace the sum by a factor of $N_{s}$}, which is the number of satellites. We define $\chi(f_{\rm gw})$ to be the angle-averaged version of the angle-dependent part of $|\mathcal{T}_{i}(f)|^{2}$:
[…]
Next, we can choose a threshold SNR value $\rho_{th}$ to invert Eq. 22 and solve for $h_{0}$. Finally, since we are specializing to monochromatic binary inspirals, we can convert this result into characteristic strain using $h_{c}\approx\sqrt{Tf_{\rm gw}}h_{0}$ Moore et al. (2015b):
[…]
@@ IV Artificial Precision Timing Array
To construct a GW detector based on PTA principles, we envision distributing periodically-flashing electromagnetic sources in the sky. These light sources need to have their periodic pulses timed to high precision so that GWs may be observed with them. This can be achieved with an array of satellites carrying atomic clocks or comparably accurate time references, which would play the same role as the astrophysical pulsars in regular PTAs. They could orbit around the Earth or the Sun. Alternatively, rather than Earth-based radio receivers, one could imagine a near-Earth receiving base to avoid atmospheric interference, as illustrated in Fig. 1. The details of this communication mechanism, including the electromagnetic pulse features and its use at detection, are beyond the scope of this paper, which focuses on clock requirements. To maintain the clock as the primary source of detector noise, the communication noise must be kept lower than the clock’s noise.
@@
Figure 1: Artist’s cartoon illustration of an APTA concept. The detector is composed of an array of satellites communicating with a depicted central receiving station. The satellites, equipped with atomic clocks or comparably accurate time references to allow for precision timing, play the role of the pulsars in this pulsar-timing-like GW detector. Objects are not drawn to scale, and the satellite array’s orbit could encircle some solar system planets or the Sun, contrary to what the cartoon figure suggests.
@@
The results of the model from Section III.2 do not depend on the pulsar/satellite communication frequency because $f_{photon}\gg f_{\rm gw}$ for reasonable signal frequency choices Isaacson (1968); Stappers et al. (2011). To observe 0.1100.1–10 Hz GWs, radio waves could do, but visible light would give a pulse-to-GW frequency ratio comparable to that of regular PTAs. X-rays may also become a good choice given recent advances in the generation of high-frequency phase-stable pulses (see e.g. Zhang et al. (2022)). We leave the optimal photon wavelength for APTA’s electromagnetic signals as a subject for future investigation.
IV.1 Detector configuration: clock precision and pulsing frequency
@@
PTAs observing for a total time $T$ with a cadence of $1/\delta t$ are generally sensitive to GW frequencies in the range $(1/T)\lesssim f_{\rm gw}\lesssim(1/\delta t)$ Moore et al. (2015b). Therefore, accessing the 0.1100.1–10 Hz band can be achieved straightforwardly by searching for GW transients and continuous signals that last at least $\sim 10$ s and taking data at least every 0.1 s.
The characteristic strain sensitivity depends linearly on the error in the PTA timing residuals $\sigma$ as indicated in Eq. 25. We will estimate the timing errors needed to conduct significant astrophysics. As we will argue, the sensitivity required is comparable to present and future atomic clock precision. We defer the discussion of additional sources of error to Section V.
@@
The best atomic clocks at present can reach fractional frequency uncertainties $\xi=\Delta f_{\rm clock}/f_{\rm clock}\sim 10^{-18}$ in 1 s of averaging and $\sim 10^{-19}–10^{-20}$ in 1 h Bothwell et al. (2022); Oelker et al. (2019); Campbell et al. (2017). In the next few years, we may reach $\xi\sim 10^{-20}$ in 1 s averaging with current 3D optical lattice clock technology. In a decade, use of squeezing and increased atom number may allow $\xi\sim 10^{-21}$ Ludlow et al. (2015). We also remark that intense efforts are underway to miniaturize and ruggedize the technology of strontium optical lattice clocks Poli et al. (2014); Koller et al. (2017); Origlia et al. (2018); Takamoto et al. (2020). Progress has been fast, which makes the prospect of deploying precision optical lattice clocks within the next decade realistic. Note that $\xi\propto t_{\rm avg}^{-1/2}$, for $t_{\rm avg}$ the time of averaging. Our averaging time is limited by the cadence, so $\xi_{\delta t}$, the relative clock uncertainty over an averaging time of $\delta t$, is
@@
where $\xi_{1\text{ s}}$ is the fractional uncertainty in 1 s of averaging. We assume APTA’s clocks reach a stable optimal fractional uncertainty $\xi_{\delta t}$ after averaging and use this value to compute the timing error as $\sigma=\xi_{\delta t}\delta t$.
@@
APTA can be sensitive to much higher frequencies than regular PTAs because its precision timing can be carried out continuously, whereas PTAs are limited by radio telescope access and, ultimately, by the folding time. Continued observation can be achieved if the beacon satellites are locked to a central observing station, and the folding time may be considerably lowered by the artificial production of a clearly modulated signal. In this case, the cadence $1/\delta t$ is bounded by the frequency at which communication pulses can be stably emitted and received.
For communication purposes, an extremely efficient way to transfer the timing precision of optical atomic clocks to the radiofrequency domain is provided by optical frequency comb technology Udem et al. (2002); Diddams et al. (2020). It has the potential to generate precision APTA pulses at a carrier frequency of 10 MHz to 100 GHz. The current state of the art for phase-stable X-ray pulses is $10^{-15}$ s Zhang et al. (2022). Moreover, the periodic light signal that enables precision timing need not come from a pulsing source. For example, high-precision spinning laser sources, cosmic lighthouses, whose low-divergence beams sweep the central station, could do. For the analysis presented, regardless of what specific mechanism produces the time-reference signal, what matters most is its ability to maintain the stability of the signal’s phase and periodicity. The form of the timing signal transmitted can range from a sinusoid to a pulse train and beyond. All of these details will be relevant in future work, but, for this study, all we need is the assumption that a moderate value of $1/\delta t=10$ kHz can be achieved with some of those technology options. Hence, we are focusing on clock uncertainty requirements by utilizing state-of-the-art low values of $\xi$ alongside conservative values of $\delta t$. The same sensitivities could be reached with higher $\xi$ by lowering $\delta t$. The determination of the optimal combination of parameters for constructing APTA is left for subsequent studies.
We return to the discussion of the assumption $\omega_{\rm gw}\tau\ll 1$ for APTA (See Section III.1). For APTA, with a clearly modulated signal, there should be no folding-like processes as required for pulsar signals, i.e. there should be no processing time considerably differentiating the pulsing frequency from the observation cadence. Therefore, the inverse cadence $\delta t$ should coincide with $\tau$. Targeting GWs in the 0.1–10 Hz frequency range, current technology is more than enough to ensure $\omega_{\rm gw}\delta t\ll 1$. Again, whether smaller values of $\delta t$ would be the optimal choice in practice is beyond the scope of this paper, but we work with Eq. 13 since it is completely realistic for APTA with respect to technological and astrophysical target requirements.
IV.2 APTA sensitivity curves